# Measuring the Mean Plane of the Distant Kuiper Belt

### v1

@@ 1 Introduction
In this Letter, we introduce a new method for measuring the mean plane of the Kuiper Belt and perform mean plane measurements at semimajor axes of $50-400\mbox{\;AU}$. In Section 2, we review the two existing methods for measuring the mean plane of a set of orbits, introduce our new method, and compare the performance of the three using simulated observations. In Section 3, we measure the mean plane for non-resonant objects in the semimajor axis bins $50-80\mbox{\;AU}$, $80-200\mbox{\;AU}$, $200-400\mbox{\;AU}$, as well as for the combined $80-400\mbox{\;AU}$ bin. We confirm the findings of previous work regarding the $50-80\mbox{\;AU}$ semimajor axis bin, and report a significant deviation from the invariable plane in the $80-200\mbox{\;AU}$ and $80-400\mbox{\;AU}$ bins. In Section 4, we investigate whether these results are consistent with unseen planets of various masses and orbits. Finally, in Section 5, we summarize our key findings and discuss implications of this work. Throughout this Letter, we refer to small bodies in the solar system with $30\mbox{\;AU}<a<1000\mbox{\;AU}$ and $q>30\mbox{\;AU}$ as trans-Neptunian objects (TNOs), but they are also commonly referred to as Kuiper belt objects (KBOs). All quoted distances are barycentric.
@@ 2 Measuring the mean plane
$L=\prod_{i=1}^{N}\frac{w(\mathbf{r}_{i})\,g(E_{i},J_{i})\,f(\hat{\mathbf{J}}_{i}\!\cdot\!\hat{\mathbf{m}})}{d\hat{\mathbf{J}}\;w(\mathbf{r}_{i})\,g(E_{i},J_{i})\,f(\hat{\mathbf{J}}\!\cdot\!\hat{\mathbf{m}})}\;\;$, (9)
@@
$p(\hat{\mathbf{J}}\cdot\hat{\mathbf{m}})=\exp\bigl{(}\gamma\,\hat{\mathbf{J}}\cdot\hat{\mathbf{m}}\bigr{)}$; (11)
@@
In order to verify that this method is truly bias-free (assuming that distribution of planes follows the von Mises distribution) and to compare the relative performance of this method to the Brown & Pan (2004) and naïve orbit-normal-averaging methods, we employ a Monte Carlo simulation involving synthetic TNO populations and hypothetical survey footprints.
@@ 4 An Unseen Planet?
Figure 3: Simulated mean plane measurement and probability ellipses for the same TNOs in Figure 2 after randomizing the $\Omega$ values and integrating for $500\mathrm{\;Myr}$ with the giant planets and an additional planet with mass $m_{p}=1\mathrm{\;M_{\oplus}}$, semimajor axis $a_{p}=160\mbox{\;AU}$, perihelion $q_{p}=150\mbox{\;AU}$, and inclination $i_{p}=20^{\circ}$. As in Figure 2, the red ‘X’ represents the invariable plane of the known planets, the blue plus marks the measurement of the mean plane using the new method described in Section 2, and the green plus marks the measurement performed with the Brown & Pan (2004) method. The purple star in each panel represents the $(p_{0},q_{0})$ location of the orbital plane of the additional planet in the simulation.
@@ Acknowledgements
We thank Renu Malhotra for helpful comments that improved the quality of the manuscript. We are pleased to acknowledge that the work reported in this paper was substantially performed using the Princeton Research Computing resources at Princeton University which is consortium of groups led by the Princeton Institute for Computational Science and Engineering (PICSciE) and Office of Information Technology’s Research Computing.

### v3

@@ 1 Introduction
In this Letter, we introduce a new method for measuring the mean plane of the Kuiper Belt and perform mean plane measurements at semimajor axes of $50-400\mbox{\;AU}$. In Section 2, we review the two existing methods for measuring the mean plane of a set of orbits, introduce our new method, and compare the performance of the three using simulated observations. In Section 3, we measure the mean plane for non-resonant objects in the semimajor axis bins $50-80\mbox{\;AU}$, $80-200\mbox{\;AU}$, $200-400\mbox{\;AU}$, as well as for the combined $80-400\mbox{\;AU}$ bin. We confirm the findings of previous work regarding the $50-80\mbox{\;AU}$ semimajor axis bin and report a significant deviation from the invariable plane in the $80-200\mbox{\;AU}$ and $80-400\mbox{\;AU}$ bins. In Section 4, we investigate whether these results are consistent with unseen planets of various masses and orbits. Finally, in Section 5, we summarize our key findings and discuss implications of this work. Throughout this Letter, we refer to small bodies in the solar system with $30\mbox{\;AU}<a<1000\mbox{\;AU}$ and $q>30\mbox{\;AU}$ as trans-Neptunian objects (TNOs), but they are also commonly referred to as Kuiper belt objects (KBOs). All quoted distances are barycentric.
@@ 2 Measuring the mean plane
where $g$ and $f$ are probability distributions. The function $g$ is spherically symmetric and $f$ incorporates the angular dependencies associated with the orientation of the disc.
@@
$L\;=\;\prod_{i=1}^{N}\frac{w(\mathbf{r}_{i})\,g(E_{i},J_{i})\,f(\hat{\mathbf{J}}_{i}\!\cdot\!\hat{\mathbf{m}})}{\displaystyle\int d\hat{\mathbf{J}}\;w(\mathbf{r}_{i})\,g(E_{i},J_{i})\,f(\hat{\mathbf{J}}\!\cdot\!\hat{\mathbf{m}})}$. (9)
@@
$f(\hat{\mathbf{J}}\cdot\hat{\mathbf{m}})=\exp\bigl{(}\gamma\,\hat{\mathbf{J}}\cdot\hat{\mathbf{m}}\bigr{)}$; (11)
@@
In order to verify that this method is truly bias-free (assuming that the distribution of planes follows the von Mises distribution) and to compare the relative performance of this method to the Brown & Pan (2004) and naïve orbit-normal-averaging methods, we employ a Monte Carlo simulation involving synthetic TNO populations and hypothetical survey footprints.
@@ 4 An Unseen Planet?
Figure 3: Simulated mean plane measurement and probability ellipses for the same TNOs in Figure 2 after randomizing the $\Omega$ values and integrating for $500\mathrm{\;Myr}$ with the giant planets and an additional planet with mass $m_{p}=0.1\mathrm{\;M_{\oplus}}$, semimajor axis $a_{p}=130\mbox{\;AU}$, perihelion $q_{p}=120\mbox{\;AU}$, and inclination $i_{p}=10^{\circ}$. As in Figure 2, the red ‘X’ represents the invariable plane of the known planets, the blue plus marks the measurement of the mean plane using the new method described in Section 2, and the green plus marks the measurement performed with the Brown & Pan (2004) method. The purple star in each panel represents the $(p_{0},q_{0})$ location of the orbital plane of the additional planet in the simulation. This is an illustrative example only.
@@ Acknowledgements
We thank Renu Malhotra and Valentin Thoss for helpful comments that improved the quality of the manuscript. We are pleased to acknowledge that the work reported in this paper was substantially performed using the Princeton Research Computing resources at Princeton University which is consortium of groups led by the Princeton Institute for Computational Science and Engineering (PICSciE) and Office of Information Technology’s Research Computing.